OpenStudy (anonymous):

Factor 10x2 - 5x A. 5x(2 - x) B. 5x(2x - 1) C. 5( 2x2 + x) D. 2(5x2 + x)

OpenStudy (nirmalnema):

b option is correct

OpenStudy (nirmalnema):

u can take out 5x common so u will get 5x(2x-1)

OpenStudy (nirmalnema):

@chulie_isobel did you understand

OpenStudy (anonymous):

kinda.... is it still the same tho? bc i forgot to raise the 2 up

OpenStudy (nirmalnema):

yaa 2x*5x=10x^2 like 2*2 = (2)^2

OpenStudy (anonymous):

so its still b?

OpenStudy (nirmalnema):

yup..

OpenStudy (anonymous):

do u think u could help me with some more problems?

OpenStudy (nirmalnema):

yup sure i will try

OpenStudy (anonymous):

thanks.. okay so heres the next oone.... Factor out the common monomial: 33x5 – 6x4 + 18x3 A. 3x3 (11x2 + 2x + 6) B. 3x3 (11x2 - 2x - 6) C. 3x3 (11x3 - 2x + 6) D. 3x3 (11x2 - 2x + 6)

OpenStudy (texaschic101):

33x^5 - 6x^4 + 18x^3 3x^3 (11x^2 - 2x + 6)

OpenStudy (nirmalnema):

D option is correct

OpenStudy (nirmalnema):

\[3x^3 (11x^2 - 2x + 6)\]

OpenStudy (anonymous):

thanks :).. my next one is: Factor: z^2 - 2z A. z(z + 2) B. (z - 1)(z - 2) C. z^2 D. z(z - 2)

OpenStudy (texaschic101):

factor out the z

OpenStudy (nirmalnema):

D again

OpenStudy (nirmalnema):

z(z-2) so u get z^2-2z

OpenStudy (anonymous):

so can u explain to me hpw to do this? i dont get it at all. the whole factoring thang...

OpenStudy (nirmalnema):

take out z (common term) from both the terms of binomial

OpenStudy (nirmalnema):

@chulie_isobel do u get it

OpenStudy (anonymous):

ya thx :3

OpenStudy (nirmalnema):

your welcome